OpenStudy (ineedserioushelp):

The midpoint of CD is E (-1,0). One endpoint is C(5,2) What are the coordinates of the other endpoint?

OpenStudy (ineedserioushelp):

I got 3,0 is that right?

OpenStudy (anonymous):

-1+x/2=5 -1+x=10 x=11 0+y/2=2 y=4 so the point is (11,4)

OpenStudy (wolfe8):

Sketching the point might help you to determine if you should add or subtract the length between midpoint and one end(if you use that method instead of the formal method as shown by @madrockz)

OpenStudy (anonymous):

ok then u show...

OpenStudy (wolfe8):

Show? You mean the informal method? I just find the difference in x and difference in y, then add it to the midpoint in the opposite direction. :) If you're doing multiple choice it's ok but for when you have to write your workings, your method is better I think.

OpenStudy (wolfe8):

My bad, I didn't check the answer. I believe the answer is (-7,-2)

OpenStudy (wolfe8):

 And then I add those differences to the midpoint in the left direction(meaning I add a negative of that difference. So basically I subtracted the midpoint with the difference)